# Unsupervised Domain Adaptation for Keyphrase Generation using Citation Contexts

### v1

@@ 1 Introduction
However, keyphrase-labeled data is notably scarce even for resource-rich languages. To date, there are only a handful of available datasets large enough to train keyphrase generation models, therefore restricting their applicability to specific domains Ye and Wang (2018); Wu et al. (2022); Garg et al. (2023). Here, we are concerned with generating keyphrases from scientific papers, for which datasets only exist in the broader scope of computer science Meng et al. (2017); Mahata et al. (2022); Garg et al. (2022) and biomedicine Houbre et al. (2022). This data scarcity issue is all the more important since current models demonstrate very limited generalization capabilities Gallina et al. (2020); Meng et al. (2021). All of this, coupled with the high computational cost of training models, underscores the necessity of developing domain adaptation methods for keyphrase generation.
@@ 3.1 Natural Language Processing (nlp)
For the nlp domain, we use the ACL Anthology Sentence Corpus [^4] https://kmcs.nii.ac.jp/resource/AASC/ that contains the sentences of 65 662 papers from the ACL Anthology up until 2022. For quality reasons, we only consider sentences from papers published in the last 20 years (2003 and upwards) and occurring within the introduction and related work sections. From these, we extracted 260 324 citation contexts with the restriction that they include at least one citation to a paper within the ACL Anthology. For each cited paper, we applied our method to extract silver-standard keyphrases from citation contexts, resulting in a confidence-ordered list of 6 199 synthetic samples. As most papers in the ACL Anthology do not provide keyphrases, we mainly relied on related conferences and journals to compile the test data for the nlp domain. More precisely, we manually collected a set of 212 documents (title and abstract) with author-assigned keyphrases from a variety of sources (e.g. LREC, SIGIR, CIKM).
@@ 3.2 Astrophysics (astro)
For the astro domain, we use the unarXive 2022 dataset Saier et al. (2023) that contains 1.9M full-text papers from arXiv. We selected the subset of 198 349 papers that belong to the Astrophysics category (astro-ph), and extracted 133 320 citation contexts originating from the introduction sections of these papers. Applying our method for each cited paper produces in a confidence-ordered list of 2 680 synthetic samples. For the astro test data, we manually collected a set of 255 documents (title and abstract) paired with author-assigned [^5] It should be noted that papers indexed with controlled vocabularies are not considered in our study. keyphrases from both arXiv and journals. To ensure topic diversity, we uniformly selected 20 documents from each astrophysics sub-category in arXiv and retrieved documents from broader-scope journals.
@@ 3.3 Paleontology (paleo)
To the best of our knowledge, there is no dataset of scientific papers available for the paleo domain. Thus, we collected 12 353 open- or free-access papers in PDF format from a wide range of journals in Paleontology. [^6] See Table 13 in Appendix A for the detailed sources. We use GROBID [^7] https://github.com/kermitt2/grobid for extracting the full-text from PDF papers, detecting inline citations and parsing bibliography, as it was shown to outperform other freely available tools Meuschke et al. (2023); Rohatgi et al. (2023). From the XML output of GROBID, we extracted 53 133 citation contexts from the introductory parts of the papers (i.e. “Introduction”, “Materials and Methods’’ and “Geological Settings”). With such a small collection, applying our method yields too few synthetic samples. To generate sufficient data for fine-tuning keyphrase generation models, we adjusted the threshold for candidate relevance (i.e. $\lambda_{r}=0.75\rightarrow 0.60$) and queried the Semantic Scholar API [^8] https://www.semanticscholar.org/ to include cited papers not present in our collection. These modifications resulted in our method generating a confidence-ordered list of 2 806 synthetic samples. For the paleo test data, we manually collected a set of 244 documents, each paired with author-assigned keyphrases, sourced from approximately 10 different journals that encompass a wide spectrum of palaeontological topics.
@@ 3.4 Statistics and Analysis
To analyze the disparities between the domains we selected, and also how they depart from KP20k (initial domain) and from other existing test datasets for keyphrase generation, we compare the main statistics of their test splits in Table 2. Here, we include three additional datasets, Inspec Hulth (2003), NUS Nguyen and Kan (2007) and SemEval-2010 Kim et al. (2010), that are composed of scientific abstracts in the computer science domain. Together with KP20k, these are probably the most commonly-used datasets for evaluating keyphrase generation models. Overall, we observe many similarities between KP20k and the test data we collected for each domain, whether in terms of the number of gold keyphrases ($\approx$5 per document), their average length ($\approx$2 tokens) or the ratio of absent keyphrases ($\approx$40%). This suggests a uniform trend in author-assigned keyphrases across scientific domains, which should facilitate generalization for keyphrase generation models. It should be noted that higher number of gold keyphrases in NUS, SemEval-2010 and Inspec stems from their distinct annotation processes, with the former two combining author- and reader-assigned keyphrases and the latter relying on professional indexers. Comparing the sizes of our domain-specific test data with those of the test splits in existing datasets shows that they are on a similar scale.
@@
Table 2: Statistics for the test data we collected for each domain in comparison with commonly used test sets.
@@ 4.1 Initial Model
We use BART Lewis et al. (2020) as our initial pre-trained language model and perform fine-tuning on the KP20k training set for 15 epochs, following Wu et al. (2023). BART was shown to yield state-of-the-art performance in keyphrase generation Zhao et al. (2022); Meng et al. (2023), surpassing other pre-trained language models, such as T5 Wu et al. (2023). Following previous work, we fine-tune BART in a One2Many setting (Yuan et al., 2020), that is, given a source text as input, the task is to generate keyphrases as a single sequence of delimiter-separated phrases. During fine-tuning, gold keyphrases are arranged in the present-absent order which was found to give the best results Meng et al. (2021). At test time, we use either greedy decoding and let the model generate the most probable keyphrases, or beam search (K=20) and assemble the top-$k$ keyphrases from all the beams as the model output.
@@ 4.3 Baselines
Although the focus of this work is domain adaption, we also provide the results of several baselines as a point of reference. The first baseline is MultiPartiteRank Boudin (2018), an unsupervised method for keyphrase extraction that leverages graph-based ranking and topical information. The second baseline is Yake Campos et al. (2020), another unsupervised method for keyphrase extraction that relies on statistical text features. The third baseline is KeyBART Kulkarni et al. (2022) in a zero-shot setting, a task-specific language model trained to learn rich representations of keyphrases. The fourth baseline is One2Set Ye et al. (2021), a Transformer-based model that uses learned control codes to generate a set of keyphrases. Trained on KP20k, this model achieves strong performance, often on-par with state-of-the-art models Wu et al. (2023).
@@ 4.5 Performance of Models on KP20k
Table 3: Performance comparison on the KP20k test set, with $\dagger$ indicating statistical significance. Scores for present and absent keyphrases separately are reported.
@@ 5.1 Confidence Ranking of synthetic samples
Table 5: Performance of BART-FT fine-tuned on the top-1K, bottom-1K and random-1K (averaged over 5 runs) samples. $\dagger$ indicates significance over BART-FT.
@@ 5.4 Bias towards Highly Cited Papers
Since our method leverages citation contexts, it produces synthetic samples that are inherently biased towards highly cited papers and their corresponding keyphrases. To investigate whether this bias is reflected in the adapted BART-FT models, we measure how often they generate keyphrases found in the synthetic samples and compare this number to that of our initial model. Results are presented in Table 8. We observe only minor differences in the number of generated keyphrases from the synthetic samples, suggesting no apparent bias. Conversely, we notice that the adapted models produce fewer of these keyphrases, as evidenced by the negative scores. We attribute this to the few-shot fine-tuning, which may not sufficiently affect the model weights to propagate bias, and reinforces the extractive capabilities of the models, thereby making them less sensitive to bias.
@@ 6 Conclusion and Future Work
In this paper, we propose silk, an unsupervised method that relies on citation contexts to create silver-standard data for adapting keyphrase generation models to new domains. We conduct experiments across three distinct scientific domains and demonstrate the effectiveness of our method for domain adaptation by few-shot fine-tuning a pre-trained model for keyphrase generation. We further validate the quality of the synthetic samples created by silk through human evaluation and analysis.

### v2

@@ 1 Introduction
However, keyphrase-labeled data is notably scarce even for resource-rich languages. To date, there are only a handful of available datasets large enough to train keyphrase generation models, therefore restricting their applicability to specific domains Ye and Wang (2018); Wu et al. (2022); Garg et al. (2023). Here, we are concerned with generating keyphrases from scientific papers, for which datasets only exist in the broader scope of computer science Meng et al. (2017); Mahata et al. (2022) and biomedicine Houbre et al. (2022). This data scarcity issue is all the more important since current models demonstrate very limited generalization capabilities Gallina et al. (2019, 2020); Meng et al. (2021). All of this, coupled with the high computational cost of training models, underscores the necessity of developing domain adaptation methods for keyphrase generation.
@@ 3.1 Natural Language Processing (nlp)
For the nlp domain, we use the ACL Anthology Sentence Corpus [^4] https://kmcs.nii.ac.jp/resource/AASC/ that contains the sentences of 65 662 papers from the ACL Anthology up until 2022. For quality reasons, we only consider sentences from papers published in the last 20 years (2003 and upwards) and occurring within the introduction and related work sections. From these, we extracted 260 324 citation contexts with the restriction that they include at least one citation to a paper within the ACL Anthology. For each cited paper, we applied our method to extract silver-standard keyphrases from citation contexts, resulting in a confidence-ordered list of 6 199 synthetic samples.
As most papers in the ACL Anthology do not provide keyphrases, we mainly relied on NLP-related conferences and journals to compile the test data for the nlp domain. More precisely, we manually collected a set of 212 documents (title and abstract) with author-assigned keyphrases from a variety of sources (e.g. LREC, SIGIR, CIKM).
@@ 3.2 Astrophysics (astro)
For the astro domain, we use the unarXive 2022 dataset Saier et al. (2023) that contains 1.9M full-text papers from arXiv. We selected the subset of 198 349 papers that belong to the Astrophysics category (astro-ph), and extracted 133 320 citation contexts originating from the introduction sections of these papers. Applying our method for each cited paper produces in a confidence-ordered list of 2 680 synthetic samples.
For the astro test data, we manually collected a set of 255 documents (title and abstract) paired with author-assigned [^5] It should be noted that controlled vocabularies are also used to index papers in astrophysics, but these are not considered in our study. keyphrases from both arXiv and journals. To ensure topic diversity, we uniformly selected 20 documents from each astrophysics sub-category in arXiv and retrieved documents from broader-scope journals.
@@ 3.3 Paleontology (paleo)
To the best of our knowledge, there is no dataset of scientific papers available for the paleo domain. Thus, we collected 12 353 open- or free-access papers in PDF format from a wide range of journals in Paleontology. [^6] See Table 13 in Appendix A for the detailed sources. We use GROBID [^7] https://github.com/kermitt2/grobid for extracting the full-text from PDF papers, detecting inline citations and parsing bibliography, as it was shown to outperform other freely available tools Meuschke et al. (2023); Rohatgi et al. (2023). From the XML output of GROBID, we extracted 53 133 citation contexts from the introductory parts of the papers (i.e. “Introduction”, “Materials and Methods’’ and “Geological Settings”). With such a small collection, applying our method yields too few synthetic samples. To generate sufficient data for fine-tuning keyphrase generation models, we adjusted the threshold for candidate relevance (i.e. $\lambda_{r}=0.75\rightarrow 0.60$) and queried the Semantic Scholar API [^8] https://www.semanticscholar.org/ to include cited papers not present in our collection. These modifications resulted in our method generating a confidence-ordered list of 2 806 synthetic samples.
For the paleo test data, we manually collected a set of 244 documents, each paired with author-assigned keyphrases, sourced from approximately 10 different journals that encompass a wide spectrum of palaeontological topics (e.g. palaeogeography, palaeoecology or stratigraphy).
@@ 3.4 Statistics and Analysis
To analyze the disparities between the domains we selected, and also how they depart from KP20k (initial domain) and from other existing test datasets for keyphrase generation, we compare the main statistics of their test splits in Table 2. Here, we include three additional datasets, Inspec Hulth (2003), NUS Nguyen and Kan (2007) and SemEval-2010 Kim et al. (2010), that are composed of scientific abstracts in the computer science domain. Together with KP20k, these are likely the most commonly-used datasets for evaluating keyphrase generation models. Overall, we observe many similarities between KP20k and the test data we collected for each domain, whether in terms of the number of gold keyphrases ($\approx$5 per document), their average length ($\approx$2 tokens) or the ratio of absent keyphrases ($\approx$40%). This suggests a uniform trend in author-assigned keyphrases across scientific domains, which should facilitate generalization for keyphrase generation models. It should be noted that higher number of gold keyphrases in NUS, SemEval-2010 and Inspec stems from their distinct annotation processes, with the former two combining author- and reader-assigned keyphrases and the latter relying on professional indexers. Comparing the sizes of our domain-specific test data with those of the test splits in existing datasets shows that they are on a similar scale.
@@
Table 2: Statistics for the test data we collected for each domain in comparison with the commonly used test sets for keyphrase generation.
@@ 4.1 Initial Model
We use BART Lewis et al. (2020) as our initial pre-trained language model and perform fine-tuning on the KP20k training set for 15 epochs, following Wu et al. (2023). BART was shown to yield state-of-the-art performance in keyphrase generation Zhao et al. (2022); Wu et al. (2022); Meng et al. (2023), surpassing other pre-trained language models, such as T5 Wu et al. (2023). Following previous work, we fine-tune BART in a One2Many setting (Yuan et al., 2020), that is, given a source text as input, the task is to generate keyphrases as a single sequence of delimiter-separated phrases. During fine-tuning, gold keyphrases are arranged in the present-absent order which was found to give the best results Meng et al. (2021). At test time, we use either greedy decoding and let the model generate the most probable keyphrases, or beam search (K=20) and assemble the top-$k$ keyphrases from all the beams as the model output. Implementation details and training times are provided in Appendix A.2.
@@ 4.3 Baselines
Although the focus of this work is domain adaption, we also provide the results of several baselines as a point of reference. The first baseline is MultiPartiteRank Boudin (2018), an unsupervised method for keyphrase extraction that leverages graph-based ranking and topical information. Despite being limited to present keyphrases, MultiPartiteRank yields the best results among non deep learning methods Do et al. (2023). We use the author’s implementation provided by the pke toolkit Boudin (2016). [^10] https://github.com/boudinfl/pke The second baseline is Yake Campos et al. (2020), another unsupervised method for keyphrase extraction that relies on statistical text features. We use the author’s implementation. [^11] https://github.com/LIAAD/yake The third baseline is KeyBART Kulkarni et al. (2022) in a zero-shot setting, a task-specific language model trained to learn rich representations of keyphrases. We use the model weights released by the authors. [^12] https://huggingface.co/bloomberg/KeyBART The fourth baseline is One2Set Ye et al. (2021), a Transformer-based model that uses learned control codes to generate a set of keyphrases. Trained on KP20k, this model achieves strong performance, often on-par with state-of-the-art models Wu et al. (2023). We use the model weights released by the authors. [^13] https://github.com/jiacheng-ye/kg_one2set
@@ 4.5 Performance of Models on KP20k
Table 3: Performance comparison of our fine-tuned BART model and baseline models on the KP20k test set, with $\dagger$ indicating statistical significance. Scores for present and absent keyphrases separately are reported.
@@ 5.1 Confidence Ranking of synthetic samples
Table 5: Performance of BART-FT fine-tuned on the top-1K, bottom-1K and random-1K (averaged over 5 runs with different seed values) samples. $\dagger$ indicates significance over BART-FT.
@@ 5.4 Bias towards Highly Cited Papers
Since our method leverages citation contexts, it produces synthetic samples that are inherently biased towards highly cited papers and their corresponding keyphrases. To investigate whether this bias is present in the adapted BART-FT models, we measure how frequently they generate keyphrases found in the synthetic samples and compare this number to that of our initial model. Results are presented in Table 8. We observe only minor differences in the number of generated keyphrases from the synthetic samples, suggesting no apparent bias. Conversely, we notice that the adapted models produce fewer of these keyphrases, as evidenced by the negative scores. We attribute this to the few-shot fine-tuning, which may not sufficiently affect the model weights to propagate bias, and reinforces the extractive capabilities of the models, thereby making them less sensitive to bias.
@@ 6 Conclusion and Future Work
In this paper, we propose silk, an unsupervised method that relies on citation contexts to create silver-standard data for adapting keyphrase generation models to new domains. We conduct experiments across three distinct scientific domains and demonstrate the effectiveness of our method for domain adaptation by few-shot fine-tuning a pre-trained model for keyphrase generation. Our results show significant improvements in in-domain performance with 1K synthetic samples over strong baselines and self-supervised domain adaptation. We further validate the quality of the synthetic samples created by silk through human evaluation and analysis.
@@
Acknowledgements
This work was supported by the French National Research Agency (ANR) through the DELICES project (ANR-19-CE38-0005-01), and by the Defense Innovation Agency (AID) and the National Centre for Scientific Research (CNRS) through the NaviTerm project (convention 2022 65 0079 CNRS Occitanie Ouest).